16. A parallel plate capacitor has dimensions 39 cm 43 cm. The plates are separated by a l 15-mm thickness of paper (dielectric constant K 3.7) What is the charge that can be stored on this capacitor, when connected to a l 4-V battery? (885 10-12 c2Nm2
Answer #1

here,

the area if each plate, A = 3.9 cm * 4.3 cm = 0.039 m * 0.043 m

thickness , d = 1.15 mm = 0.0015 m

k = 3.7

V = 1.4 V

the charge stored in the capacitor, Q = C * V

Q = ( K * area * e0 /d) * V

Q = ( 3.7 * 0.039 * 0.043 * 8.85 * 10^-12 /0.0015) * 1.4 C

Q = 5.13 * 10^-11 C

the charge stored in the capacitor is 5.13 * 10^-11 C